OpenStudy (yolo729):

The mass of a freight train is 9.98 ⋅ 10^6 kilograms. The mass of an aircraft carrier is 7.3 ⋅ 10^7 kilograms. Approximately how many more kilograms is the mass of an aircraft carrier than the mass of a freight train? Show your work and write your answer in scientific notation.

OpenStudy (yolo729):

@3mar

OpenStudy (3mar):

Thank you.

OpenStudy (yolo729):

welcome

OpenStudy (3mar):

I think it is a very basic simple subtraction

OpenStudy (3mar):

Can you carry on?

OpenStudy (yolo729):

730,000,000-998,000,000

OpenStudy (3mar):

where is the step?

OpenStudy (yolo729):

it ends up being a negative

OpenStudy (yolo729):

or did i do it wrong

OpenStudy (marihelenh):

It shouldn't be negative.

OpenStudy (yolo729):

what did i do wrong

OpenStudy (phi):

you should do 7.3 ⋅ 10^7 - 9.98 ⋅ 10^6 notice the exponents are different. you should change 7.3 ⋅ 10^7 to 73 ⋅ 10^6 (notice we multiplied 7.3 by 10 and to compensate, reduced 10^7 to 10^6) now you have 73 ⋅ 10^6 - 9.98 ⋅ 10^6

OpenStudy (3mar):

The mass of an aircraft carrier - the mass of a freight train=7.3(10^7)-9.98(10^6)=63,020,000=63.02(10^6)

OpenStudy (marihelenh):

To find what 9.98 x 10^6 kg is, you move the decimal place over six places to the right. For 7.3 x 10^7 you move the decimal place over seven spots to the right.